OpenStudy (denise):

How do I write slope intercept form of the equation Through(2,1) slope =-3/2

OpenStudy (anonymous):

To find the equation of a line given a point \[(x _{0},y _{0})\], you can use the formula: \[y-y_{0} = m(x-x _{0})\] where m = the slope of the line. So you just sub in your point and you slope and solve for y.

OpenStudy (denise):

thanks! I have lots of various questions

OpenStudy (anonymous):

An alternate way is to substitute your given point and slope into the equation y=mx+b and solve for b, and then just rewrite the equation in the proper form.